# SVD-Free Low-Rank Adaptive Gradient Optimization for Large Language Models

### v1

@@ Abstract
Low-rank optimization has emerged as a promising direction in training large language models (LLMs) to reduce the memory usage of adaptive optimizers by constraining learning to a lower-dimensional space. Prior work typically projects gradients of linear layers using approaches based on Singular Value Decomposition (SVD). However, applying SVD-based procedures individually to each layer in large models is computationally expensive and incurs additional memory costs due to storing the projection matrices. In this work, we propose a computationally efficient and conceptually simple two-step procedure to approximate SVD-based gradient projections into lower-dimensional spaces. First, we construct a complete orthogonal basis using predefined orthogonal matrices of the Discrete Cosine Transform (DCT). Second, we adaptively select basis columns based on their alignment with the gradient of each layer. Each projection matrix in our method is obtained via a single matrix multiplication followed by a lightweight sorting step to identify the most relevant basis vectors. Due to the predefined nature of the orthogonal bases, they are computed once at the start of training. During training, we store only the indices of the selected columns, avoiding the need to store full projection matrices for each layer. Our numerical experiments on both pre-training and fine-tuning tasks demonstrate the effectiveness of our dual strategy in approximating optimal low-rank projections, matching the performance of costly SVD-based methods while achieving faster runtime and reduced memory usage.
@@ 1 Introduction
The Adam optimizer (Kingma and Ba, 2014) and its regularized version AdamW (Loshchilov and Hutter, 2019) have become the standard approach for optimizing deep neural networks in various settings. With the recent increase in scale of LLMs up to billions and trillions of parameters, training with AdamW becomes more and more challenging, as its internal state requires two momentum buffers that scale with the model size. This practical problem paved the way for a line of research that focuses on reducing the memory usage of optimizer states in the context of adaptive gradient optimization. These approaches range from quantizing the states to 8 bits (Dettmers et al., 2021) to the recent GaLore optimizer (Zhao et al., 2024) inspired from the LoRA techniques (Hu et al., 2021; Lialin et al., 2023), that compresses the gradient matrix using low-rank decomposition based on SVD. Several improvements to GaLore have been proposed to enhance its performance, such as LDAdam (Robert et al., 2025), FRUGAL (Zmushko et al., 2024), FIRA (Chen et al., 2024), BAdam (Luo et al., 2024), Q-GaLore (Zhang et al., 2024b). The key aspect of GaLore and its later improvements is the low-rank decomposition based on matrix factorization, such as SVD or QR decomposition. These techniques are known to be computationally intensive as they have to be invoked for each linear layer, either at each step (for, e.g., LDAdam) or once at a few steps (for, e.g., GaLore). To mitigate the high cost of SVD, we ask: can we find an alternative projection to serve as an accurate replacement for the orthogonal matrices in SVD, that is much cheaper to compute, and portable across memory-efficient optimizers?
@@
Contributions. In this work, we address our research question by proposing a cheaper alternative to the orthogonal matrices computed via SVD. Concretely, we propose using the orthogonal matrices from the general class of Discrete Fourier Transforms, such as the Discrete Cosine Transform (DCT), which has been successfully used in image compression for the JPEG algorithm. To the best of our knowledge, we are the first to use it in the context of low-rank adaptive gradient optimization. We summarize our contributions as follows:
  • We show the DCT matrix is an accurate replacement for the orthogonal matrices computed via SVD. We reduce the running time and improve the memory usage as we store only one DCT matrix for the entire network, computed once at the beginning of the training. Instead of storing one projection matrix for each layer, we just have to store $r$ (rank) integers, representing the indices of most significant columns for that particular layer.
@@
  • We propose a theory-based approach to compute the most significant columns of the DCT matrix to obtain a dynamic projection matrix tailored to each layer.
  • In order to show that DCT matrices accurately approximate the orthogonal matrices of SVD, we add the DCT transformation into existing low-rank optimizers and show they recover training loss, memory usage and/or running time.
@@
  • We propose a standalone optimizer that uses the DCT projection and optionally adds quantized error feedback for the projection error, rotates the momentum buffers such that new low-rank gradients are correctly incorporated and allows changing the low-rank subspace every step.
2 Related Work
We review the prior work in the literature focused on reducing the running time and memory usage of optimizers by compressing the gradient, especially by low-rank compression. Most approaches use individual projection matrices tailored to the gradient at each layer by invoking techniques based on quantization or matrix factorization, such as SVD, QR or PCA. One pioneering work in the context of low-rank adaptive optimization is the recent GaLore optimizer (Zhao et al., 2024) which performs SVD once at a few steps to project the gradient to a lower-dimensional space. GaLore was followed by several other optimizers that improve certain aspects of it.
The first improvement aspect is the running time, which forces GaLore to compute SVD once at a few steps in order to be feasible in practice for large models. LDAdam (Robert et al., 2025) compresses the first order momentum in AdamW, replaces SVD with a block power-iteration (Bentbib and Kanber, 2015) and performs a smooth subspace transition by rotating the first and second momentum accordingly to incorporate gradients from the same subspace at each step.
The second improvement aspect is convergence. By default, GaLore discards the projection error, which LDAdam stores and incorporates into the gradient at the next step. In contrast, the recent work FRUGAL (Zmushko et al., 2024) makes the distinction between the low-rank gradient (called state-full) and the projection error (called state-free). The state-full gradient is used in AdamW, while the state-free gradient is fed to an optimizer without state, such as SignSGD. The main idea is to leverage the remaining gradient information in the projection error since it is available at each step instead of discarding or storing it. A concurrent work to FRUGAL is FIRA (Chen et al., 2024), which aims to preserve the low-rank constraint for memory efficiency while achieving full-rank performance by proper scaling of the projection error.
In addition to improvements over GaLore, some prior work focuses on achieving memory usage similar to SGD while preserving the performance of AdamW. Apollo (Zhu et al., 2024) exploits the redundancy in the computed adaptive learning rates of AdamW and shows that element-wise learning rate is not necessary. As a consequence, they focus on computing one adaptive learning rate per channel, which significantly improves the memory usage and throughput by allowing larger batch sizes due to memory savings. They achieve this by projecting the gradient to a lower-dimensional space using a random matrix only to compute the scaling for the gradient, which remains full-rank.
Another approach worth mentioning is Online Subspace Descent Liang et al. (2024), which replaces SVD projection with Online PCA, which involves computing the projection matrix as a solution of an optimization function focused on 1) minimizing the projection error and 2) forcing the projection matrix to be orthogonal. The authors indicate that performing one step with Adam to solve this additional optimization problem is enough to obtain a qualitatively good projection matrix $P$. In contrast, our method does not introduce such overheads during training since the DCT projection is already computed at the beginning of training.
As complementary approaches to low-rank compression, there is some prior work worth mentioning due to their similar goal of reducing the memory usage and running time with a different methods than low-rank. AdamW-8bit (Dettmers et al., 2021) quantizes the states of AdamW optimizer to 8-bits using efficient CUDA kernels, with the effect of reducing the memory usage by a half compared to bfloat16 and without introducing additional computational overheads. In the same spectra of methods we find MicroAdam (Modoranu et al., 2024), where the optimizer states are compressed via sparsity and error feedback. Adam-mini (Zhang et al., 2024a) reduces optimizer states memory by assigning a single learning rate to each parameter block carefully partitioning the parameter matrix into blocks.
Despite all aforementioned approaches being similar in using SVD-based projections, they differ in a few aspects. The most important one in our view is how they handle the projection error and how often they update the low-dimensional subspace, which we clarify in Table 1.
Our approach uses DCT projection coupled with a dynamic column selection to determine a projection matrix tailored to the gradient for a particular layer, can use any update frequency and can be integrated into any optimizer, regardless of the way it handles the projection error.
Low-rank Projection Type Frequency* Error
GaLore (Zhao et al., 2024) SVD 200 discard
Fira (Chen et al., 2024) SVD 200 norm-based scaling
Subspace Descent (Liang et al., 2024) Online PCA 1 discard
LDAdam (Robert et al., 2025) Block Power Iteration 1 error feedback
DCT-AdamW (this work) DCT 1 error feedback
@@ 3 Method
This section introduces the orthogonal DCT matrix we use to replace SVD, the technique to dynamically choose columns to obtain a projection matrix tailored to the gradient at each layer, as well as the standalone optimizer DCT-AdamW we propose.
@@ 3.1 Discrete Cosine Transform
The Discrete Cosine Transform (DCT) is widely used in the signal processing and data compression (e.g., JPEG algorithm for image compression) and consists of $n$ orthogonal basis vectors whose components are cosines (Strang, 1999). There are minor variations of DCT and in this work we will use DCT-2 and DCT-3 (one is the transpose of the other). We denote the DCT-3 matrix of order $n$ by $Q\in\mathbb{R}^{n\times n}$, defined as $Q_{ij}=\sqrt{2/n}\cdot\cos\frac{i(2j+1)\pi}{2n}$, with $i,j\in[n]$, where the first row has to be divided by $\sqrt{2}$ in order for $Q$ to be orthogonal, i.e. $Q^{\top}Q=I_{n}$. We can efficiently compute $Q$ using a vectorized implementation that is fast for large values of $n$ on the GPU as follows. We first create one vector $L=[0,\cdots,n-1]^{\top}$, which is used to create the matrix $\mathcal{I}\in\mathbb{N}^{n\times n}$ by replicating $L$ on the columns of $\mathcal{I}$:
We restate that we need to divide the first row by $\sqrt{2}$ to make $Q$ orthogonal. The computational efficiency comes from the element-wise product $\mathcal{I}\odot(2\mathcal{I}^{\top}+1)$ that actually computes the integer entries $i(2j+1)$. Once we compute $Q$, we can use it as a replacement for the SVD procedure in conjunction with the dynamic selection of columns, which we explain in detail next.
3.2 Dynamic Selection of Columns
General View. After computing the DCT matrix $Q\in\mathbb{R}^{n\times n}$, we need to compute a projection matrix $Q_{r}\in\mathbb{R}^{n\times r}$ by selecting $r$ columns from $Q$ to project the gradient $G\in\mathbb{R}^{n\times n}$ to an $r$-dimensional space $g=GQ_{r}\in\mathbb{R}^{n\times r}$. First, we compute $S=GQ$ containing the scalar products between rows of $G$ and columns of $Q$. We rank the columns of $S$ based on their $L_{1}$-norm and then pick the largest $r$ columns, which we use in $Q_{r}$. The $i^{th}$ column in matrix $S$ contains the scalar products between rows of $G$ and column $i$ of $Q$. We view these scalar products as similarities and we want to choose the columns of $Q$ with largest similarity with rows of $G$. Since the columns of $Q$ are defined as cosines of different frequencies at each component, we state that columns of $Q$ act as a filter for the rows of $G$. We choose $L_{1}$-norm because we do not want the similarities to be altered, as it would be the case for $L_{2}$-norm. Using this approach, we ensure a dynamic mechanism to choose the most appropriate columns of $Q$ for the current gradient matrix $G$. Each layer will have its own set of $r$ indices for the columns. The case illustrated so far is for a general square matrix $G$. In the context of GaLore, where $G\in\mathbb{R}^{n\times m}$, we reduce the smallest dimension of $G$ to $r$, which means we have to take into consideration two cases, based on the relationship between $n$ and $m$.
The first case is $n\geq m$, which is known as “right projection” in GaLore, where $Q\in\mathbb{R}^{m\times m}$ is the DCT-3 matrix. We compute $S=GQ$ and the corresponding $Q_{r}\in\mathbb{R}^{m\times r}$ contains the most significant $r$ columns of $S$. The second case is $n<m$, which is known as “left projection” and this time we compute $S=Q^{\top}G$, where we are interested in the rows of $Q$ and columns of $G$. Here, $Q^{\top}\in\mathbb{R}^{n\times n}$ is the DCT-2 matrix (the transpose of DCT-3, but we store only the DCT-3 matrix, $Q$) and in order to choose the columns of $Q$ we have to select $r$ indices of the largest rows in $S$.
Usually, the smallest dimension of the gradient matrix is the hidden size of the model $d_{\rm model}$. As a result, the memory overhead of replacing the SVD projection with the DCT matrix and the dynamic selection of columns is storing a single matrix $Q\in\mathbb{R}^{d_{\rm model}\times d_{\rm model}}$ for the entire model and $r$ integers for the corresponding columns from $Q$ for each layer.

### v2

# FFT-based Dynamic Subspace Selection for Low-Rank Adaptive Optimization of Large Language Models

@@ Abstract
Low-rank optimization has emerged as a promising direction in training large language models (LLMs) to improve running time and reduce the memory usage of adaptive optimizers by constraining learning to a lower-dimensional space. Prior work typically projects gradients of linear layers using approaches based on Singular Value Decomposition (SVD) or QR-decomposition. Applying these techniques individually to each layer in large models is computationally expensive and incurs additional memory costs due to storing the projection matrices. In this work, we propose a computationally efficient and conceptually simple, two-step procedure to approximate SVD/QR-based gradient projections into lower-dimensional spaces by using a predefined orthogonal matrix of the Discrete Cosine Transform (DCT). We dynamically select columns from the DCT matrix based on their alignment with the gradient of each layer. The effective projection matrices are obtained via a simple matmul with the DCT matrix in $O(n^{3})$ time, followed by a lightweight sorting step to identify the most relevant basis vectors. For large layers, DCT can be computed via Makhoul’s $N$-point algorithm based on Fast Fourier Transform (FFT) in $O(n^{2}\log(n))$ time. Due to the predefined nature of the orthogonal bases, they are computed once at the start of training. Our numerical experiments on both pre-training and fine-tuning tasks demonstrate the effectiveness of our dual strategy in approximating optimal low-rank projections, obtaining an approach with rank-independent runtime that matches the performance of costly SVD/QR-based methods while achieving faster runtime and reduced memory usage by up to $25\%$ across different model sizes.
@@ 1 Introduction
The Adam optimizer (Kingma & Ba, 2014) and its regularized version AdamW (Loshchilov & Hutter, 2019) have become the standard approach for optimizing deep neural networks in various settings. With the recent increase in scale of LLMs up to billions and trillions of parameters, training with AdamW becomes more and more challenging, as its internal state requires two momentum buffers that scale with the model size. This practical problem paved the way for a line of research that focuses on reducing the memory usage of optimizer states in the context of adaptive gradient optimization. These approaches range from quantizing the states to 8 bits (Dettmers et al., 2021) to the recent GaLore optimizer (Zhao et al., 2024) inspired from the LoRA techniques (Hu et al., 2021; Lialin et al., 2023), that compresses the gradient matrix using low-rank decomposition based on SVD. Several improvements to GaLore have been proposed to enhance its performance, such as LDAdam (Robert et al., 2025), FRUGAL (Zmushko et al., 2024), FIRA (Chen et al., 2024), BAdam (Luo et al., 2024), Q-GaLore (Zhang et al., 2024). The key aspect of GaLore and its later improvements is the low-rank decomposition based on matrix factorization, such as SVD or QR decomposition.
Recently, the Muon (Jordan et al., 2024) optimizer sparked the community’s attention due to the faster convergence in pretraining settings, achieved by orthogonalizing the momentum matrix using Newton-Schulz Newton-Schulz, an iterative procedure difficult to parallelize for large-scale pretraining runs as the full-sized matrices must be materialized on GPU. The Dion optimizer (Ahn et al., 2025) aims to reduce this overhead by employing low-rank, orthogonal updates. However, it requires storing a projection matrix for each layer and uses QR-decomposition to orthogonalize the low-rank components, making its running time dependent on the rank.
All these SVD/QR-based techniques are known to be computationally intensive as they have to be invoked for each linear layer, either at each step (for, e.g., LDAdam, Muon) or once at a few steps (for, e.g., GaLore). To mitigate the high computational and memory costs of these procedures, we ask: can we find an alternative low-rank projection approach to serve as an accurate replacement for the orthogonal matrices in SVD/QR for low-rank compression of optimizer states, that is much cheaper to compute, and portable across memory-efficient optimizers?
@@
Contributions. In this work, we address our research question by proposing a cheaper alternative to the orthogonal matrices computed via SVD/QR. Concretely, we propose using the orthogonal matrices from the general class of Discrete Fourier Transforms, such as the Discrete Cosine Transform (DCT), which has been successfully used in image compression for the JPEG algorithm. To the best of our knowledge, we are the first to use it in the context of low-rank adaptive gradient optimization. We summarize our contributions as follows:
  • We propose a dynamic column selection approach to adaptively choose columns from a fixed orthogonal matrix to compute a low-rank projection of the input matrix. The effective projection matrix is obtained from the fixed orthogonal matrix by indexing the columns (Section 2.1);
  • We motivate that DCT is a good candidate for our dynamic column selection approach due to the reduced time complexity to compute the alignments with the input matrix via the Makhoul’s $N$-point algorithm to compute a fast DCT with $O(n^{2}\log(n))$ time complexity compared to $O(n^{3})$ incurred by a basic matrix multiplication (Makhoul, 1980). This can yield speedups of up to $8-50\times$ in computing the alignments for large layers (Appendix C).
@@
  • We theoretically justify our dynamic column selection approach to compute the most significant columns of any fixed orthogonal matrix (including the DCT-based) to obtain a projection matrix tailored to each layer (Section 4);
  • We show the DCT-based projection is a fast and accurate replacement for the orthogonal matrices computed via SVD/QR for the Adam-like optimizers (such as FRUGAL, FIRA, LDAdamW) in the context of low-rank compression of optimizer states. We reduce the running time and improve the memory usage as we store only one DCT matrix per GPU for the entire network, computed once at the beginning of the training. In addition, we store only $r$ (rank) integers, representing the indices of the most significant columns for each layer instead of storing one projection matrix.
  • We propose two standalone optimizers that use the DCT-based dynamic column selection approach:
    1. 1.
      Trion (Section 2.3), which improves Dion by replacing the Power-Iteration with our DCT-based dynamic column selection approach to compute a low-rank representation of the momentum buffer, followed by orthogonalization via Newton-Schulz iteration. We provide a DDP-compatible implementation that communicates orthogonal, low-rank momentum across devices and compute the final layer updates locally using the projection matrix obtained from DCT. To the best of our knowledge, we are the first ones to reduce the complexity of Newton-Schulz using a low-rank approximation of momentum.
    2. 2.
@@
      DCT-AdamW (Section 2.4), which replaces the SVD-based low-rank projections and optionally adds quantized error feedback for the projection error, rotates the momentum buffers such that new low-rank gradients are correctly incorporated and allows changing the low-rank subspace every step;
@@ 2 Method
This section presents our dynamic column selection approach that works with any orthogonal matrix, briefly introduces the Discrete Cosine Transform (DCT) and the motivation of using it and finally introduce the two algorithms we propose: Trion as improvement to Dion to replace Power-Iteration and DCT-AdamW as an improvement to low-rank AdamW variants to replace SVD/QR.
2.1 Dynamic Column Selection
General View. Given an orthogonal matrix $Q\in\mathbb{R}^{n\times n}$, we need to compute a projection matrix $Q_{r}\in\mathbb{R}^{n\times r}$ by selecting $r$ columns from $Q$ to project the gradient $G\in\mathbb{R}^{n\times n}$ to an $r$-dimensional space $g=GQ_{r}\in\mathbb{R}^{n\times r}$. First, we compute the similarities matrix $S=GQ$ containing the scalar products between rows of $G$ and columns of $Q$. We rank the columns of $S$ based on their $\ell_{1}$- or $\ell_{2}$-norm and then pick the indices of the largest $r$ columns according to the chosen norm, which we use to index columns in $Q$ to obtain $Q_{r}$. The $i^{th}$ column in matrix $S$ contains the scalar products between each row of $G$ and $i^{th}$ column of $Q$, as detailed in Appendix B. We view these scalar products as similarities (or alignments) of rows in $G$ with columns of $Q$ and we want to choose the columns of $Q$ with largest similarity with rows of $G$. Using this approach, we ensure a dynamic mechanism to choose the most appropriate columns of $Q$ for the current gradient matrix $G$ to minimize the projection error (see Section 4). Each layer will have its own set of $r$ indices for the columns. The dynamic behavior comes from the $n$-choose-$r$ possible sets of indices to select from.
The rule of thumb in the low-rank factorization methods for optimization we are targeting in this work is to compress the smallest dimension of a matrix to $r$ dimensions (e.g. from $\mathbb{R}^{n\times m}$ to $\mathbb{R}^{n\times r}$ for $n\geq m$). Usually, the smallest dimension of the gradient matrix is $d_{\rm model}$, the hidden (embedding) size of the model. As a result, the memory overhead of our dynamic column selection approach is the cost of storing only one orthogonal matrix $Q\in\mathbb{R}^{d_{\rm model}\times d_{\rm model}}$ (DCT in our case) per GPU for the entire model and $r$ integers for the corresponding column indices from $Q$ for each layer to create $Q_{r}$.
@@ 2.2 Discrete Cosine Transform
The Discrete Cosine Transform (DCT) is widely used in the signal processing and data compression literature (e.g., JPEG algorithm for image compression) and consists of $n$ orthogonal basis vectors whose components are cosines (Strang, 1999). There are minor variations of DCT and in this work we will use DCT-II/III. We denote the DCT-III matrix of order $n$ by $Q\in\mathbb{R}^{n\times n}$, defined as $Q_{ij}=\sqrt{2/n}\cdot\cos\frac{i(2j+1)\pi}{2n}$, with $i,j\in[n]$, where the first row has to be divided by $\sqrt{2}$ in order for $Q$ to be orthogonal, i.e. $Q^{\top}Q=I_{n}$. The DCT-II matrix is the transpose of DCT-III. In Appendix A, we provide details on how we can efficiently materialize the DCT matrix on GPU. Once we compute $Q$, we can use it in the dynamic column selection approach to efficiently compute low-rank projections of any two-dimensional matrix.
In Appendix C, we discuss why we choose DCT matrix. In short, the particular structure of DCT allows us to enable fast computation for the similarities matrix $S$ in $O(n^{2}\log(n))$ time using the Makhoul’s $N$-point algorithm (see Appendix D) instead of $O(n^{3})$ time for basic matmul.
2.3 Trion: DCT-based improvement to Dion
In this section, we present how we can apply the DCT-based dynamic column selection approach to compute low-rank projection of momentum in the context of Dion (Ahn et al., 2025) to obtain a faster and more accurate optimizer, which we call Trion.
Dion uses Power-Iteration to compute a low-rank projection and orthogonalizes it using QR-decomposition, whose running time depends on the rank $r$. Instead, we propose replacing these techniques with a rank independent approach to compute the low-rank projection based on DCT matrix, which requires computing the similarity matrix $S_{t}$, followed by a top-$r$ ranking to determine the index of columns that best align with the momentum matrix, denoted by the set $i_{t}$. The matrix $S_{t}$ represents the DCT of the momentum matrix and it can be computed using the Makhoul’s algorithm or simply by only one matrix multiplication $S_{t}=B_{t}D_{C}$, where $B_{t}\in\mathbb{R}^{R\times C}$ is the momentum and $D_{C}\in\mathbb{R}^{C\times C}$ is the DCT matrix.
Once we determine the indices $i_{t}$ of the most significant columns in $D_{C}$, we can extract the low-rank momentum $b_{t}$ from the similarity matrix $S_{t}$, as well as the projection matrix $Q_{t}$ by indexing $D_{C}$, which will be used in (a) computing the projection error $\Delta_{t}$ and (b) projecting the orthogonal low-rank momentum back to the higher dimensional space to update the model.
We would like to emphasize that we input the low-rank momentum $b_{t}\in\mathbb{R}^{R\times r}$ to Newton-Schulz and not the original momentum buffer $B_{t}\in\mathbb{R}^{R\times C}$, which significantly reduces the FLOPs. Moreover, we can use the efficient triton kernels provided in the Muon implementation from the official Dion repository111github.com/microsoft/dion to speed up the computations even further for large ranks $r$.
Communication in Distributed Training. We base the Trion implementation on the published code for the Muon 222github.com/KellerJordan/Muon/blob/master/muon.py#L138 optimizer that leverages the ZeRO (Rajbhandari et al., 2020) approach in data-parallel settings. Specifically, the model update $O_{t}$ for a layer is computed only on one GPU, and the result is communicated to other GPUs using all-gather, drastically reducing the computation costs for large models. Since we replicate the DCT matrix on each GPU, we would like to emphasize that we communicate only low-rank terms $o_{t}$ from the source GPU to other devices and perform the step $O_{t}=o_{t}Q_{t}^{\top}$ locally on each device. This way, we do not communicate full size, orthogonal matrices $O_{t}$, but only low-rank versions $o_{t}$, thus reducing the overall iteration time.
This is particularly useful in the FSDP-compatible implementation, where we can materialize the full low-rank matrices $o_{t}$ on each device using all-to-all (see this Muon implementation 333github.com/microsoft/dion/blob/main/dion/muon.py#L22) to perform Newton-Schulz on a low-rank input, followed by a resharding. After resharding, each GPU can compute the individual slice of $O_{t}$ and update its own shard. However, the FSDP implementation requires deciding how each layer should be sharded based on whether that particular layer requires a left- or right-projection in order to avoid unwanted calls to communication primitives to materialize full tensors to compute the alignments $S$.
Algorithm 1 Trion Optimizer
1:Define the DCT-II/III matrix $D_{C}\in\mathbb{R}^{C\times C}$
2:for $t=\{1,2,\dots,T\}$ do
3:  $G_{t}=\nabla_{\theta}L(\theta_{t})\in\mathbb{R}^{R\times C}$
4:  $B_{t}=M_{t-1}+G_{t}\in\mathbb{R}^{R\times C}$
5:  $S_{t}=\textsc{Makhoul}(B_{t})\in\mathbb{R}^{R\times C}$ $\triangleright$ or $S_{t}=B_{t}\cdot D_{C}$ to compute similarities;
6:  $i_{t}=\textsc{DynamicColumnSelection}(S,r)\in\mathbb{N}^{r}$ $\triangleright$ select largest $r$ columns by $\ell_{1}/\ell_{2}$-norm
7:  $Q_{t}=D_{C}[:,i_{t}]\in\mathbb{R}^{C\times r}$ $\triangleright$ projection matrix containing most significant $r$ columns
8:  $b_{t}=S_{t}[:,i_{t}]\in\mathbb{R}^{R\times r}$ $\triangleright$ extract low-rank momentum from the similarity matrix $S$
9:  $\Delta_{t}=B_{t}-b_{t}Q_{t}^{\top}$ $\triangleright$ the error is $b_{t}Q_{t}^{\top}$ and replaces $P_{t}R_{t}^{\top}$ error from Dion
10:  $M_{t}=\mu B_{t}+(1-\mu)\Delta_{t}=B_{t}-(1-\mu)b_{t}Q_{t}^{\top}$ $\triangleright$ update momentum with error feedback
11:  $o_{t}=\textsc{NewtonSchulz}(b_{t})\in\mathbb{R}^{R\times r}$ $\triangleright$ Newton-Schulz applied on low-rank $b_{t}$, can use triton kernels
12:  $O_{t}=o_{t}Q_{t}^{\top}\in\mathbb{R}^{R\times C}$ $\triangleright$ project orthogonalized low-rank momentum to original size
13:  $\theta_{t+1}=(1-\lambda\eta_{t})\theta_{t}-\eta_{t}\;\max(1,\sqrt{R/C})\;O_{t}\in\mathbb{R}^{R\times C}$
14:end for